OpenStudy (anonymous):

1 byte = 2^3 bits. How many bits is this in simplified form? How many bits is 5 bytes? 1 kilobyte (kB) is 2^10B. How many bytes is this in simplified form? __B How many bits is this? __b How many bits is 30kB? __b 1 megabyte(MB) = 2^10 kB. How many kB is this in simplified form? __kB How many bytes is this? __B 1 gigibyte(GB)=2^10 MB. How many bytes is this in simplified form? _B How many bits are there in a GB? _b How many MB in 2GB? How many Mb is this?

Parth (parthkohli):

What is \(2^3\)? What is \(2^{10 \times B}\)? What is \(30 \times 8\)? What is \(2^{10}\)? What is one byte in KB, and one byte in MB?

Parth (parthkohli):

Try your homework before posting...

OpenStudy (anonymous):

I'll try and make it a little bit simpler, 1byte is 8 bits, 1 kilobyte is 1024 bytes, 1 megabyte is 1024 kilobytes, 1 gigabyte is 1024 megabytes, i know it's not listed but 1 terabyte is 1024 gigabytes and so on...

OpenStudy (anonymous):

so the first one would be 20?

OpenStudy (anonymous):

1.How many bits is 5 bytes? Answer of this would be 40?

OpenStudy (anonymous):

yes, 5x8 = 40

OpenStudy (anonymous):

2. 1 byte = 2^3 bits. How many bits is this in simplified form? this owuld be 6 right?

OpenStudy (anonymous):

2^3 = 2x2x2 = 8

OpenStudy (anonymous):

3,1 kilobyte (kB) is 2^10B. How many bytes is this in simplified form? __B How many bits is this? __b im lost on this one i know that its 20 but i know the rest